OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 6, using the expression -2(5)n - 1. 1,223 -1,023 7,812 -7,812

OpenStudy (anonymous):

@dan815

OpenStudy (welshfella):

Sn = a1 * r^n - 1 -------- r - 1

OpenStudy (welshfella):

here a1 = -2 , r = 5, n = 6

OpenStudy (anonymous):

Sn = -2 * 5^6 - 1 = 15624 -------- ------- = -7812.5 5 - 1 4

OpenStudy (anonymous):

Is that right?

OpenStudy (welshfella):

-7812

OpenStudy (anonymous):

Can you help me with this one: Find the sum of a finite arithmetic sequence from n = 1 to n = 15, using the expression 2n + 5.

OpenStudy (welshfella):

i'm assuming 2n + 5 is the formula for the nth term so first term would be 2+ 5 = 7 and 15th term would be 2(15) + 5 = 35 Sn = (n/2) ( a + l) here n = number of terms, a = first term and l = last ( 15th) term

OpenStudy (anonymous):

I dont understand what to do